How to solve by elimination?
x+y=7
x+3y=11

Answer by stanbon(75887)

How to solve by elimination?
x+y=7
x+3y=11
-----------------
Subtract and solve for "y":
2y = 4
y = 2
-----
Solve for "x":
x + y = 7
x + 2 = 7
x = 5
================
Cheers,
Stan H.
==============